OpenStudy (anonymous):

a cylinder has a surface area of 339 cm squared. The radius is two times greater than the height. What is the radius of the cylinder? answer to the nearest cm.

OpenStudy (anonymous):

You have two pieces of information: r = 2h (the radius is twice the height) A = 339 cm^2 The formula for surface area of a cylinder is: A = (2 * pi * r^2) + (2 * pi * r * h) Express h in terms of r: h = r / 2 Substitute that for h: 339 = (2 * pi * r^2) + (2 * pi * r * r/2) Multiply out the second parentheses: 339 = (2 * pi * r^2) + (pi * r^2) Simplify: 339 = 3 * pi * r^2 Divide both sides by 3 * pi: r^2 = 339 / (3 * pi) Simplify: r^2 = 113 / pi = 36 (rounded) r = sqrt(36) = 6 cm

OpenStudy (anonymous):

okay. Thanks so much! But where are you getting the 2 from, cause i know its 2 times greater, but still, I dont understand this. can you do a step by step, showing what to do....

OpenStudy (anonymous):

That is step by step! hehe :) Medal? :D

OpenStudy (anonymous):

yeah, but how would I write that on paper?

OpenStudy (dan815):

lol future u are so funny!!!

OpenStudy (anonymous):

haha but I dont get it!!